OpenStudy (anonymous):

Is my answer right? What is the solution to the rational equation 3/x + 3 = 2/x - 2? x = 12

hero (hero):

\[\frac{3}{x+3} = \frac{2}{x-2}\] 2(x + 3) = 3(x - 2) 2x + 6 = 3x - 6 6 + 6 = 3x - 2x 12 = x Correct

OpenStudy (anonymous):

Is that 3/(x+3) = 2/(x-2) or (3/x) + 3 = (2/x) - 2 ?

OpenStudy (anonymous):

thank you. xo